# Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation

### v1

@@ 4.2 The ORBIT trials
Another pair of clinical trials that investigated the effect of ciprofloxacin are the ORBIT 3 and ORBIT 4 trials [40]. The trials assessed the effect of inhaled liposomal ciprofloxacin compared to placebo in patients with non-cystic fibrosis bronchiectasis and chronic lung infection with Pseudomonas aeruginosa. Like the RESPIRE trials, the ORBIT trials also showed considerable heterogeneity. Figure 4 shows p-values, point estimates, and CIs for the primary endpoint (time to the first exacerbation; effect quantified with a log hazard ratio) and a secondary endpoint (frequency of exacerbations; effect quantified with a log rate ratio). Table 3 gives numerical summaries.
Figure 4: Results of the ORBIT trials [40] for the effect of ciprofloxacin in patients with non-cystic fibrosis bronchiectasis and chronic lung infection with Pseudomonas aeruginosa. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods are shown along with corresponding median estimates and CIs (95% and 99.875% via telescope lines). All p-values are based on the alternative ”greater” and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
@@
Table 3: Point estimates, 95% CIs, and p-values for the ORBIT trials [40].
Lower CI Estimate Lower CI CI width P-value (one-sided)
@@
   ORBIT 3 $-0.34$ $-0.01$ $0.32$ $0.66$ $0.47636$
     ORBIT 4 $-0.63$ $-0.33$ $-0.03$ $0.60$ $0.01657$
   Two-trials rule $-0.39$ $-0.10$ $0.16$ $0.55$ $0.22692$
     Meta-analysis $-0.41$ $-0.18$ $0.04$ $0.45$ $0.05305$
   Tippett $-0.48$ $-0.24$ $0.02$ $0.50$ $0.03286$
     Fisher $-0.45$ $-0.21$ $0.03$ $0.49$ $0.04610$
   Pearson $-0.40$ $-0.15$ $0.12$ $0.52$ $0.14328$
     Edgington $-0.45$ $-0.18$ $0.12$ $0.57$ $0.12149$
@@
   ORBIT 3 $-0.43$ $-0.16$ $0.11$ $0.54$ $0.12083$
     ORBIT 4 $-0.73$ $-0.46$ $-0.19$ $0.54$ $0.00036$
   Two-trials rule $-0.47$ $-0.24$ $-0.02$ $0.45$ $0.01460$
     Meta-analysis $-0.51$ $-0.31$ $-0.12$ $0.38$ $0.00062$
   Tippett $-0.60$ $-0.39$ $-0.16$ $0.44$ $0.00072$
     Fisher $-0.57$ $-0.35$ $-0.14$ $0.43$ $0.00048$
   Pearson $-0.49$ $-0.28$ $-0.05$ $0.44$ $0.00765$
     Edgington $-0.57$ $-0.31$ $-0.06$ $0.51$ $0.00734$
Looking at the secondary endpoint, there is also considerable heterogeneity between the results from ORBIT 3 ($\log\widehat{\mathrm{RR}}=-0.16$ with 95% CI from $-0.43$ to $0.11$) and ORBIT 4 ($\log\widehat{\mathrm{RR}}=-0.46$ with 95% CI from $-0.73$ to $-0.2$) leading to some more noticeable qualitative differences between the methods. That is, the 99.875% CIs from meta-analysis and Fisher’s method exclude a log rate ratio of zero while the remaining methods include it, leading to different decisions at the $0.025^{2}$ level. Again, Edgington’s CI is much wider than the others due to the substantial heterogeneity.
@@ 5 Extension to more than two trials
The methods discussed so far have focused on the setting where only two trials are available, but in practice it may happen that investigators want to assess the combined evidence from more than two trials. In this context, Rosenkranz [31] suggested that decision rules should maintain the type I error rate of the two-trials rule for two studies $\alpha^{2}$, even if there are more than two studies. This can be implemented using combined p-value functions, as all methods considered before can be generalized to more than two trials [13, 27]. A decision rule can then be based on the combined one-sided p-value for the null hypothesis of no effect or a $(1-2\alpha^{2})\times 100\%$ CI obtained from a combined p-value function. In addition, a point estimate and 95% CI can be used to summarize the combined evidence.
While all point estimates and confidence intervals in this setting can be computed numerically, some of the analytical results derived earlier generalize to more than two studies. Specifically, closed-form median estimates and confidence intervals remain available for the two-trials rule, Tippett’s method, and meta-analysis, whereas such closed-form solutions are not available for Fisher’s, Pearson’s, and Edgington’s methods [27]. In particular, for Edgington’s method, one might expect the median estimate (16) to generalize by incorporating additional effect estimates with inverse standard error weights. However, a comparison with numerically computed median estimates showed that this is not the case. Thus, the inverse standard error weighted average in (16) corresponds to Edgington’s median estimate only in the setting of two trials.
Figure 5: Combined p-value function based on combining all four estimates from the RESPIRE trials [6, 7, 8] along with corresponding median estimates and CIs (95% and 99.875% via telescope lines).
Figure 5 shows p-value functions that combine all four results from the two RESPIRE trials, as also done by Chotirmall and Chalmers [8] with fixed-effect meta-analysis. Looking at the median estimates, we see the same patterns as when the methods are applied to only two trials; The median estimates from meta-analysis and Edgington’s method are somewhere in between the trials’ individual estimates. The median estimates based on Fisher’s and Tippett’s method are the most anti-conservative, and the median estimates based on Pearson’s method and the two-trials rule are the most conservative. A decision rule that maintains the $\alpha^{2}=0.025^{2}$ type I error rate of the two trials rule could now be defined by flagging drug efficacy when the $(1-2\alpha^{2})\times 100\%=99.875\%$ excludes the null value. Following this rule, we can see that Fisher’s method and meta-analysis flag efficacy while the remaining methods do not. In sum, this example illustrates how combined p-value functions can be applied to more than two trials, while allowing to maintain the same type I error rate as for two trials.
@@ 6 Discussion
The two-trials rule has been widely discussed in the literature but discussions have mostly focused on hypothesis testing characteristics, such as power or type I error rate [9, 3, 4, 35, 5, 41, 29, 42, 43, 10, 30, 31, 22]. In this paper, we took a different perspective, systematically examining the two-trials rule and alternative methods in terms of effect estimation. By casting them in a combined p-value function framework, we derived compatible p-values, confidence intervals, and point estimates. These quantities are compatible in the sense that the two-sided p-value for a null value is less than $\alpha$ if and only if the null value is excluded by the $(1-\alpha)\times 100\%$ CI, and that the point estimate is contained in the CI at any confidence level. While meta-analytic effect estimates, CIs, and p-values have been well studied, our novel results enable computation of CIs and effect estimates based on the two-trials rule. Investigators could therefore report not only individual trial p-values (essentially the two-trials rule) and meta-analytic estimates but also point estimates and CIs based on the two-trials rule.
@@
Our findings also clarify how different p-value combination methods implicitly target different estimands. Reassuringly, under effect homogeneity (i.e., the same true effect in both trials), all methods yield consistent point estimates and CIs that shrink toward the true effect as standard errors decrease, although some show bias. Theoretically, meta-analysis has the smallest variance among all unbiased estimators (attaining the Cramér-Rao lower bound) and may be preferred. However, under effect heterogeneity – arguably the more realistic scenario – it is less clear which method should be recommended. The two-trials rule and Pearson’s method are conservative (targeting the less extreme effect), Fisher’s and Tippett’s methods are anti-conservative (targeting the more extreme effect), while Edgington’s method and meta-analysis are balanced (targeting a weighted average). This raises an important question: What kind of effect is of scientific interest when the true trial effects differ? If the investigators are interested in the less extreme effect – arguably a sensible choice when the effects relate to a medical treatment with potential side effects – then the two-trials rule and Pearson’s method seem reasonable. On the other hand, a weighted average effect, as targeted by meta-analysis and Edgington’s method, might be a relevant estimand if it is representative for a larger population [33]. Finally, the more extreme effect might be the relevant estimand if the maximum achievable benefit of a treatment is of scientific interest, in which case Tippett’s and Fisher’s methods might be reasonable choices.
@@
Beyond theoretical considerations, practical issues must be addressed. A major concern is that if the effect estimates from both trials are the same, the two-trials rule, Tippett’s, Fisher’s, and Pearson’s methods all produce counterintuitive effect estimates that differ from the one observed in both trials. Such point estimates are unintuitive and difficult to communicate to non-statisticians. From this perspective, Edgington’s method and meta-analysis may be preferable. In particular, Edgington’s method can also account for effect heterogeneity by widening its CI when there is heterogeneity, and it asymptotically always includes both effects. Another challenge is aligning decisions based on a one-sided combined p-value thresholded at $\alpha$ with two-sided CIs. This requires using a $(1-2\alpha^{2})\times 100\%$ confidence level. However, in many fields, researchers are not used to such confidence levels, so we suggest to report both a more conventional level (e.g., 95%) along with $(1-2\alpha^{2})\times 100\%$ via telescope-style CIs, as well as the underlying p-value function, as in Figures 35.
@@
Our results have broader implications beyond the two-trial setting. Methods for combining p-values are also used in adaptive trials, where they enable combination of stage-wise p-values [44, 45, 46]. Combined p-value functions can be generalized to more than two studies, making them applicable to meta-analysis [13, 27]. They can also be applied to replication and real-world evidence studies, where the two-trials rule (under different names such as significance criterion or vote-counting) is used to assess the replicability of original findings [47, 48, 39]. In all these scenarios, we may consider combined p-value functions for parameter estimation, but in each application researchers must also decide which combination method has the statistical properties to estimate the scientific effect of interest. Future research may also examine other combination methods beyond the ones considered here, such as the inverse chi-square method [49, 17], random-effects meta-analysis [32], and combining p-value functions that are based on the exact distribution of the data rather than normality [e.g., the p-value function based on Fisher’s exact test with mid-p correction as considered in 50, 27]. Additionally, fixed-effect meta-analysis has a Bayesian interpretation, corresponding to posterior inferences assuming equal true study effects and a flat prior distribution. Investigating whether other p-value combination methods have similar Bayesian justifications could be an interesting avenue for future work. To sum up, combined p-value functions provide a unified approach for combining results from two trials that can be further developed theoretically. Moreover, our software implementation allows researchers to conveniently apply these methods in practice.
Conflict of interest
We declare no conflict of interest.
Software and data
Data from the RESPIRE trials were extracted from Table 3 in De Soyza et al. [7] and Table 3 in Aksamit et al. [6]. Data from the ORBIT trials were extracted from page 219 in Haworth et al. [40]. Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/twotrials. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.15017483. We used the statistical programming language R version 4.4.1 (2024-06-14) for analyses [51] along with the confMeta [52], ggplot2 [53], xtable [54], kableExtra [55], dplyr [56], and knitr [57] packages.
Appendix A The twotrials R package
We have developed the twotrials R package for easily conducting combined p-value function inference based on the parameter estimates (with standard errors) from two trials. The package can be installed from GitHub via the R command remotes::install_github(repo = "SamCH93/twotrials", subdir = "package") (requires the remotes package). We plan to make the package available on the Comprehensive R Archive Network (CRAN) after receiving community feedback.
For every p-value combination method discussed in this paper, the package provides a combined p-value function (e.g., pEdgington) and a combined estimation function (e.g., muEdgington). While these can be used to manually compute p-values and parameter estimates, the convenience function twotrials automatically computes estimates and p-values based on all methods, and allows for easy printing and plotting of the results. The following code chunk illustrates its usage by reproducing the results for the 14-day treatment group from Table 2.
@@ Appendix A The twotrials R package
library(twotrials) # load package
## combine logRR estimates from RESPIRE trials
results <- twotrials(null = 0, t1 = -0.4942, t2 = -0.1847, se1 = 0.1833,
                     se2 = 0.1738, alternative = "less", level = 0.95)
print(results, digits = 2) # print summary of results
## INDIVIDUAL RESULTS
##    Trial Lower CI Estimate Upper CI P-value
##  Trial 1    -0.85    -0.49    -0.13  0.0035
##  Trial 2    -0.53    -0.18     0.16  0.1440
## COMBINED RESULTS
##           Method Lower CI Estimate Upper CI P-value
##  Two-trials rule    -0.57    -0.28   -0.011  0.0207
##    Meta-analysis    -0.58    -0.33   -0.084  0.0043
##          Tippett    -0.68    -0.39   -0.084  0.0070
##           Fisher    -0.64    -0.35   -0.087  0.0043
##          Pearson    -0.58    -0.32   -0.044  0.0114
##        Edgington    -0.64    -0.34   -0.048  0.0109
## NOTES
## Confidence level: 95%
## Null value: 0
## Alternative: less
plot(results, xlim = c(-1, 0.5), two.sided = TRUE) # plot p-value functions

### v2

@@ 4.2 The ORBIT trials
Another pair of clinical trials that investigated the effect of ciprofloxacin are the ORBIT 3 and ORBIT 4 trials 41. The trials assessed the effect of inhaled liposomal ciprofloxacin compared to placebo in patients with non-cystic fibrosis bronchiectasis and chronic lung infection with Pseudomonas aeruginosa. Like the RESPIRE trials, the ORBIT trials also showed considerable heterogeneity. Figure 4 shows p-values, point estimates, and CIs for the primary endpoint (time to the first exacerbation; effect quantified with a log hazard ratio) and a secondary endpoint (frequency of exacerbations; effect quantified with a log rate ratio). Table 3 gives numerical summaries.
Figure 4: Results of the ORBIT trials 41 for the effect of ciprofloxacin in patients with non-cystic fibrosis bronchiectasis and chronic lung infection with Pseudomonas aeruginosa. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods are shown along with corresponding median estimates and CIs (95% and 99.875% via telescope lines). All p-values are based on the alternative "greater" and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
@@
Table 3: Point estimates (with implicit weights), 95% CIs (with widths), and p-values for the ORBIT trials 41. Note that weigths and CI widths are computed from unrounded numbers and may not always exactly correspond to the rounded point estimates and CIs.
Log rate ratio Weight ORBIT 3 95% CI CI width P-value (one-sided)
@@
[…]
@@
[…]
Looking at the secondary endpoint, there is also considerable heterogeneity between the results from ORBIT 3 ($\log\widehat{\mathrm{RR}}=-0.16$ with 95% CI from $-0.43$ to $0.11$) and ORBIT 4 ($\log\widehat{\mathrm{RR}}=-0.46$ with 95% CI from $-0.73$ to $-0.19$) leading to some more noticeable qualitative differences between the methods. That is, the 99.875% CIs from meta-analysis and Fisher’s method exclude a log rate ratio of zero while the remaining methods include it, leading to different decisions at the $0.025^{2}$ level. Again, Edgington’s CI is much wider than the others due to the substantial heterogeneity.
@@ 5 Extension to more than two trials
The methods discussed so far have focused on the setting where only two trials are available, but in practice it may happen that investigators want to assess the combined evidence from more than two trials. In this context, Rosenkranz 32 suggested that decision rules should maintain the type I error rate of the two-trials rule for two studies $\alpha^{2}$, even if there are more than two studies. This can be implemented using combined p-value functions, as all methods considered before can be generalized to more than two trials 13, 28. A decision rule can then be based on the combined one-sided p-value for the null hypothesis of no effect or a $(1-2\alpha^{2})\times 100\%$ CI obtained from a combined p-value function. In addition, a point estimate and 95% CI can be used to summarize the combined evidence.
While all point estimates and CIs in this setting can be computed numerically, some of the analytical results derived earlier generalize to more than two studies. Specifically, closed-form median estimates and CIs remain available for the two-trials rule, Tippett’s method, and meta-analysis, whereas such closed-form solutions are not available for Fisher’s, Pearson’s, and Edgington’s methods 28. In particular, for Edgington’s method, one might expect the median estimate (18) to generalize by incorporating additional effect estimates with inverse standard error weights. However, a comparison with numerically computed median estimates showed that this is not the case. Thus, the inverse standard error weighted average in (18) corresponds to Edgington’s median estimate only in the setting of two trials.
Figure 5 shows p-value functions that combine all four results from the two RESPIRE trials, as also done by Chotirmall and Chalmers 8 with fixed-effect meta-analysis. Looking at the median estimates, we see the same patterns as when the methods are applied to only two trials; The median estimates from meta-analysis and Edgington’s method are somewhere in between the trials’ individual estimates. The median estimates based on Fisher’s and Tippett’s method are the most anti-conservative, and the median estimates based on Pearson’s method and the two-trials rule are the most conservative. A decision rule that maintains the $\alpha^{2}=0.025^{2}$ type I error rate of the two trials rule could now be defined by flagging drug efficacy when the $(1-2\alpha^{2})\times 100\%=99.875\%$ excludes the null value. Following this rule, we can see that Fisher’s method and meta-analysis flag efficacy while the remaining methods do not. In sum, this example illustrates how combined p-value functions can be applied to more than two trials, while allowing to maintain the same type I error rate as for two trials.
Figure 5: Combined p-value function based on all four estimates from the RESPIRE trials 6, 7, 8 along with corresponding median estimates and CIs (95% and 99.875% via telescope lines).
@@ 6 Discussion
The two-trials rule has been widely discussed in the literature but discussions have mostly focused on hypothesis testing characteristics, such as power or type I error rate 9, 3, 4, 36, 5, 42, 30, 43, 44, 10, 31, 32, 22. In this paper, we took a different perspective, systematically examining the two-trials rule and alternative methods in terms of effect estimation. By casting them in a combined p-value function framework, we derived compatible p-values, confidence intervals, and point estimates. These quantities are compatible in the sense that the two-sided p-value for a null value is less than $\alpha$ if and only if the null value is excluded by the $(1-\alpha)\times 100\%$ CI, and that the point estimate is contained in the CI at any confidence level. While meta-analytic effect estimates, CIs, and p-values have been well studied, our novel results enable computation of CIs and effect estimates based on the two-trials rule. Investigators could therefore report not only individual trial p-values (essentially the two-trials rule) and meta-analytic estimates but also point estimates and CIs based on the two-trials rule.
@@
Our findings also clarify how different p-value combination methods implicitly target different estimands. Reassuringly, under effect homogeneity (i.e., the same true effect in both trials), all methods yield consistent point estimates and CIs that shrink toward the true effect as standard errors decrease, although some show bias. Theoretically, meta-analysis has the smallest variance among all unbiased estimators (attaining the Cramér-Rao lower bound) and may be preferred. However, under effect heterogeneity – arguably the more realistic scenario – it is less clear which method should be recommended. The two-trials rule and Pearson’s method are conservative (targeting the less extreme effect), Fisher’s and Tippett’s methods are anti-conservative (targeting the more extreme effect), while Edgington’s method and meta-analysis are balanced (targeting a weighted average). This raises an important question: What kind of effect is of scientific interest when the true trial effects differ? If the investigators are interested in the less extreme effect – arguably a sensible choice when the effects relate to a medical treatment with potential side effects – then the two-trials rule and Pearson’s method seem reasonable. On the other hand, a weighted average effect, as targeted by meta-analysis and Edgington’s method, might be a relevant estimand if it is representative for a larger population 34. Finally, the more extreme effect might be the relevant estimand if the maximum achievable benefit of a treatment is of scientific interest, in which case Tippett’s and Fisher’s methods might be reasonable choices. This parallels the findings of Heard and Rubin-Delanchy 21, who showed that many p-value combination methods are equivalent to a likelihood ratio test for specific alternative hypotheses. This means that each such method can be most powerful under certain conditions. Therefore, researchers must carefully reflect which alternative hypothesis is most relevant to their application – just as they need to reflect on choosing an appropriate estimand – to select a suitable combination method.
@@
Beyond theoretical considerations, practical issues must be addressed. A major concern is that if the effect estimates from both trials are the same, the two-trials rule, Tippett’s, Fisher’s, and Pearson’s methods all produce counterintuitive effect estimates that differ from the one observed in both trials. Such point estimates are unintuitive and difficult to communicate to non-statisticians. Moreover, only Edgington’s method and meta-analysis produce the same combined estimate and two-sided confidence interval in case the alternative of the combined $p$-values is changed 28, which seems another practically desired property. From this perspective, Edgington’s method and meta-analysis may be preferable. In particular, Edgington’s method can also account for effect heterogeneity by widening its CI when there is heterogeneity and asymptotically always includes both effects. However, this is traded off with a less efficient median estimate under effect homogeneity, whose standard error can even be larger than those from both trials if they greatly differ. Finally, another practical challenge is aligning decisions based on a one-sided combined p-value thresholded at $\alpha^{2}$ with two-sided CIs. This requires using a $(1-2\alpha^{2})\times 100\%$ confidence level. However, in many fields, researchers are not used to such confidence levels, so we suggest to report both a more conventional level (e.g., 95%) along with $(1-2\alpha^{2})\times 100\%$ via telescope-style CIs, as well as the underlying p-value function, as in Figures 35. The idea of telescope-style CIs is not new but has been suggested before in different contexts 45.
A broader issue is the question of whether two trials are actually necessary. If the designs of the two trials are so similar that they can be considered exchangeable (“direct replications” 46), there are various arguments in favor of conducting one large trial instead of two smaller ones 9, 10. Also our study demonstrates that having two trials instead of one makes estimation more complicated. Conversly, if the trial designs differ significantly (“conceptual replications” 46, e.g., if they use different endpoints or populations), achieving success in both trials may provide more robust evidence of treatment efficacy. From this perspective, it is sensible to design the trials differently to some extent 10. However, there is perhaps a limit to how different the trials can be, as when there is too much heterogeneity, combining the effect estimates into a single number would no longer be meaningful.
@@
Our results have broader implications beyond the two-trial setting. Methods for combining p-values are also used in adaptive trials, where they enable combination of stage-wise p-values 47, 48, 49. Combined p-value functions can be generalized to more than two studies, making them applicable to meta-analysis 13, 28. They can also be applied to replication and real-world evidence studies, where the two-trials rule (under different names such as significance criterion or vote-counting) is used to assess the replicability of original findings 50, 51, 40. In all these scenarios, we may consider combined p-value functions for parameter estimation, but in each application researchers must also decide which combination method has the statistical properties to estimate the scientific effect of interest. Future research may also examine other combination methods beyond the ones considered here, such as the inverse chi-square method 52, 17, the harmonic mean $\chi^{2}$ test 53, the Cauchy combination test 54, random-effects meta-analysis 33, and combining p-value functions that are based on the exact distribution of the data rather than normality, e.g., the p-value function based on Fisher’s exact test with mid-p correction 55, 28. Additionally, fixed-effect meta-analysis has a Bayesian interpretation, corresponding to posterior inferences assuming equal true study effects and a flat prior distribution. Investigating whether other p-value combination methods have similar Bayesian justifications could be an interesting avenue for future work. To sum up, combined p-value functions provide a unified approach for combining results from two trials that can be further developed theoretically. Moreover, our software implementation allows researchers to conveniently apply these methods in practice.
@@
*Acknowledgments
We thank the editor, associate editor, Stephen Senn, and another anonymous reviewer for valuable comments that led to numerous additions and improvements. The acknowledgment of these individuals does not imply their endorsement of the paper.